OpenStudy (ultrasoundwave15):

Simplfy 6^1/3 x 36^1/3 36 6 sqrt6 3sqrt6

OpenStudy (ultrasoundwave15):

@I_Always_Smiling can u help me?

OpenStudy (michele_laino):

please note that, if I apply the property of multiplication of powers having the same base, I can write this: \[6^{1/3}*6^{2/3}=6^{\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }}=...\]

OpenStudy (ultrasoundwave15):

So, would I calculate that?? ^

OpenStudy (michele_laino):

what is:\[\frac{ 1 }{ 3 }+\frac{ 2 }{ 3 }=...\]

OpenStudy (ultrasoundwave15):

3/3

OpenStudy (michele_laino):

namely 1

OpenStudy (ultrasoundwave15):

ok

OpenStudy (michele_laino):

then: \[6^{1}=...\]

OpenStudy (ultrasoundwave15):

36

OpenStudy (michele_laino):

\[6^{1}=6\]

OpenStudy (ultrasoundwave15):

oops, I thought wrong

OpenStudy (ultrasoundwave15):

sorry

OpenStudy (michele_laino):

no problem

OpenStudy (michele_laino):

welcome to OpenStudy!

OpenStudy (ultrasoundwave15):

Thankyou! How'd you know I was knew?

OpenStudy (ultrasoundwave15):

new*

OpenStudy (michele_laino):

by my teaching experience

OpenStudy (ultrasoundwave15):

oh, thanks for your help!

OpenStudy (michele_laino):

thanks!